topic:- Cost accounting

* no hand writing

*the answer must be unique not copied "plagiarized "

Assignment Question(s):​​​​​​

Q1. Spaghetti is a family-owned restaurant in KSA. The corporate office provides 2 kinds of services (maintenance and personnel) to 3 locations (Riyadh, Jeddah and Dammam).

In a recent accounting period, the cost for maintenance was $25,000 and for personnel was $15,000. Maintenance costs are allocated on the basis of square feet, and personnel costs are allocated based on the basis of number of employees.

​​Number of

​Square Feet​Employees

Support Departments:

Maintenance​1,000​15

Personnel​1,000​8

Operating Departments:

Riadh​5,000​12

Jeddah​6,000​18

Dammam​3,000​10

If the restaurant uses the direct method and the costs of Riadh location, Jeddah location and Dammam location are $15,000, $12,000 and $10,000 respectively, calculate the total costs allocated to each restaurant location.

Q 2 The cost to produce one unit of the product is:

​​Material ​​​  $ 12.00

​​Labor ​​​​ $ 9.00

​​Variable cost​​​ $ 6.00

Fixed expenses​​  $ 18.00

​​​​

Total fixed expenses: $ 1,440,000

The company’s normal capacity is 100,000 units. The figures given above are for 80,000 units.

The company has received a special offer for 20,000 units for a price of $ 36 per unit from a foreign customer.

Advice the manufacturer on whether the order should be accepted.  

Q 3 Taha Compagny produces three products: Product A, Product B, and Product C.

  During the year the joint costs of processing the three products were $400,000. Production and sales value information were as follows:

​​Sales Value

Product​Units​at Split-Off​Separable Costs​Selling Price

A​400,000​   $20 per unit​   $ 20 per unit​$40 per unit

B​400,000​   $18 per unit​   $ 15 per unit​   $28 per unit

C​800,000​$12 per unit​   $14 per unit​   $17 per unit

a.​Allocate the joint costs using the physical output method.

b.​Allocate the joint costs using the net realizable value method.

c. ​Allocate the joint costs using sales value at split-off point method.
Answer #1

Ques! Riyadh Teddah Dammam 1560o 12.000 Direct costs maintainance cost 10000 8929 10714 5357 Personnel cost 4500 6750 3750 28(6) Net Realisable value selling price - separable cost A = 40-20-20 B - 28-15 - 13 = 17- 14 = 3 costs A= B: 222222 144444 33